# Multi-neutron correlations in light nuclei via ab-initio lattice simulations

### v1

@@ I.1 Uncertainty quantification of chiral forces
We start from 284 non-implausible N3LO chiral interactions on the lattice [1], obtained via history matching [2, 3, 4, 1] that explicitly account for uncertainties in NN scattering data, the truncation of the chiral hierarchy, and binding energies of selected nuclei. To construct the posterior distribution informed by experimental data, the first step typically involves incorporating additional observables associated with the relevant structural features, thereby conditionally refining the physical credibility of the non-implausible interaction set. The posterior is subsequently obtained within a Bayesian framework that combines a Gaussian likelihood with a prior over the chiral interactions,
@@ I.4 Schematic illustration of four-neutron correlations in 7H
Figure S1 schematically illustrates the angular coordinates used to characterize the four-body correlations in light nuclei, using 7H as an example. The four neutrons can be partitioned into two neutron-neutron pairs. We adopt the same two-body angular definitions for each pair as in the two-body correlation function, and define the pair vectors and as the relative displacement vectors connecting the two neutrons within each pair. The intra-pair opening angles and quantify the internal compactness of the pairs, while and encode the degree of radial asymmetry between the two neutrons within a pair (as in the two-body analysis). The angle characterizes the relative opening between the pair c.m. vectors, and the torsion angle measures the non-coplanarity (twist) between the two neutron-neutron pair vectors. This set of angles provides a convenient parameterization of the four-neutron geometry and underlies the angular distributions discussed in the main text.
@@ I.5 Extrapolation in Euclidean time
where denotes the -th eigenstate, is its energy gap to the g.s., and represents its overlap with the initial state relative to that of the g.s.. Following the procedure in Ref. [17], we first fit the energies calculated with the simple Hamiltonian using the lowest order of Eq. (S6) to extract and . These parameters are then used in a second-stage fit for general operators, such as higher-order energy correlations and observables that commute with the Hamiltonian, based on the functional form
@@ I.8 Two-neutron correlations in 6He and 3H
Figure S2(a) shows the two-neutron correlation density of the valence neutrons in 6He. Since 6He has been extensively studied both experimentally and theoretically [36, 37, 38], it provides a benchmark for our correlation-analysis procedure. Figure S2(b) shows the two-neutron correlation density of 3H for comparison with 7H. Note that the 7H correlations are normalized by the total number of neutron pairs, . Therefore, the per-pair short-distance component shown for 7H is reduced by a factor of 15 relative to the single-pair 3H benchmark purely due to this normalization convention. The two-neutron correlation density characterizes the spatial extent and angular structure of the neutron pair in 3H and exhibits a similar structure to the short-distance component of the multi-neutron correlation patterns obtained in 7H.
@@ I.10 Opening angle and dependence in four-neutron correlation of 7H
To assess potential systematic effects associated with the marginal distribution, we examine the stability of the opening-angle distribution with respect to variations of the intra-pair neutron-neutron angle selection and the intra-pair neutron-neutron opening-angle cut. Here, represents the angle between the relative position vector of the two nucleons and the bisector of their opening angle ( quantifies the deviation from radial symmetry). As shown in Fig. S4(a), changing from to at fixed leaves the normalized distribution essentially unchanged. When we vary the selection window, it primarily changes the effective statistics (i.e., the number of accepted configurations) but leaves the relative distribution essentially unchanged, indicating no discernible change in the underlying four-neutron correlation pattern within uncertainties. In this work, we select to balance the radial symmetry and statistics. In Fig. S4(b), we vary the intra-pair neutron-neutron opening-angle cut while keeping fixed. Because directly quantifies the internal geometry of an neutron-neutron pair, this cut defines which intra-pair correlation patterns are included in the sampled configurations. Guided by the two-body correlation analysis discussed above, we find that the dominant strength associated with compact, dineutron-like correlations in 7H is concentrated primarily within , which motivates our baseline choice. Accordingly, for moderate variations of the cut (–), the normalized distribution exhibits only minor differences and remains consistent within uncertainties. This indicates that our baseline selection captures the relevant neutron-neutron correlation physics without either overly truncating compact configurations or introducing spurious distortions.
@@
I.11 Internal angular and radial distributions from the four-neutron correlations in 7H
@@
Figure S5 provides additional information on the internal angular and radial distributions in the four-neutron correlation analysis and serves as a consistency check against the two-neutron correlations in 7H. Panels (a,e) compile the joint probability densities and for the two neutron-neutron pairs identified in each sampled four-neutron configuration. The bulk of the probability density is concentrated around the typical dineutron angular and radial scales extracted independently from the two-neutron correlations (dashed lines), indicating that the leading internal scales inferred from two-body observables are consistent with those obtained from the explicit four-body correlation. This agreement supports the interpretation of the observed bimodal angular pattern in 7H from the two-body correlation analysis: the two-body correlation function represents a statistical distribution over all neutron-neutron pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. Panels (b,f), obtained with the selection , are similar to (a,e) and contain the majority of the sampled configurations, indicating that the (extended) sector dominates the sampled four-body ensemble. Panels (d,h) correspond to a characteristic window around a maxima of the marginal distribution (). For the large- window (d,h), the radial distribution shows that and are comparable, and shifts toward smaller intra-pair opening angles, consistent with a near back-to-back arrangement of the two pair c.m. vectors at comparable radii and a more compact intra-pair geometry. In contrast, the small- window (c,g) shows a more asymmetric joint radial distribution, featuring enhanced weight near fm, alongside a nearly symmetric contribution at fm. The intra-pair opening-angle distribution shifts to larger angles and becomes noticeably broader, with enhanced support around , compared to the tighter concentration near in panel (d). The associated torsion-angle distribution is also weighted toward large in our calculation, indicating a strongly non-coplanar (twisted) and interlaced four-neutron geometry with pronounced four-body correlations.

### v2

@@ I.1 Uncertainty quantification of chiral forces
We start from 282 non-implausible N3LO chiral interactions on the lattice [1], obtained via history matching [2, 3, 4, 1] that explicitly account for uncertainties in NN scattering data, the truncation of the chiral hierarchy, and binding energies of selected nuclei. To construct the posterior distribution informed by experimental data, the first step typically involves incorporating additional observables associated with the relevant structural features, thereby conditionally refining the physical credibility of the non-implausible interaction set. The posterior is subsequently obtained within a Bayesian framework that combines a Gaussian likelihood with a prior over the chiral interactions,
@@ I.4 Schematic illustration of four-neutron correlations in 7H
Figure S1 schematically illustrates the angular coordinates used to characterize the four-body correlations in light nuclei, using 7H as an example. The four neutrons are grouped into two neutron-neutron pairs, and for each pair we adopt the same two-body angular variables as in the two-body correlation function. We define the pair vectors and as the relative displacement vectors connecting the two neutrons within each pair. The intra-pair opening angles and quantify the compactness of the pairs, while and encode the degree of radial asymmetry between the two neutrons within a pair (as in the two-body analysis). The angle characterizes the relative opening between the pair c.m. vectors, and the torsion angle measures the non-coplanarity (twist) between the two neutron-neutron pair vectors. This set of angles provides a convenient parameterization of the four-neutron geometry and underlies the angular distributions discussed in the main text.
@@ I.5 Extrapolation in Euclidean time
where denotes the -th eigenstate, is its energy gap to the g.s., and represents its overlap with the initial state relative to that of the g.s. Following the procedure in Ref. [17], we first fit the energies calculated with the simple Hamiltonian using the lowest order of Eq. (S6) to extract and . These parameters are then used in a second-stage fit for general operators, such as higher-order energy correlations and observables that commute with the Hamiltonian, based on the functional form
@@ I.8 Two-neutron correlations in 6He and 3H
Figure S2(a) shows the two-neutron correlation density of the valence neutrons in 6He. Since 6He has been extensively studied both experimentally and theoretically [36, 37, 38], it provides a benchmark for our correlation analysis. Figure S2(b) shows the two-neutron correlation density of 3H for comparison with 7H. Note that the 7H correlations are normalized by the total number of neutron pairs, . Therefore, the per-pair short-distance component shown for 7H is reduced by a factor of 15 relative to the 3H reference purely due to this normalization convention. The two-neutron correlation density characterizes the spatial extent and angular structure of the neutron pair in 3H and closely resembles the short-distance component of the multi-neutron correlation patterns obtained in 7H.
@@ I.10 Opening angle and dependence in four-neutron correlation of 7H
We assess the robustness of the inter-pair opening-angle distribution to variations of the intra-pair angle selections used to identify dineutron-like neutron pairs in Figs. S4. Here is the inter-pair opening angle between the two pair c.m. position vectors (measured from the system c.m.). The intra-pair geometry is characterized by the neutron–neutron opening angle and by , which quantifies the deviation from radial symmetry (where is defined as the angle between the relative position vector of the two neutrons and the bisector of their opening angle). As shown in Fig. S4(a), changing from to at fixed leaves the normalized distribution essentially unchanged. When we vary the selection window, it primarily changes the accepted sample size (i.e., the number of accepted configurations) but leaves the normalized shape of the distribution essentially unchanged, indicating no discernible change in the underlying four-neutron correlation pattern within uncertainties. In this work, we select to balance the radial symmetry and statistics. In Fig. S4(b), we vary the intra-pair neutron-neutron opening-angle cut while keeping fixed. Because directly quantifies the internal geometry of a neutron-neutron pair, this cut defines which intra-pair correlation patterns are included in the sampled configurations. Guided by the two-body correlation analysis discussed above, we find that the dominant strength associated with compact, dineutron-like correlations in 7H is concentrated primarily within , which motivates our baseline choice. Accordingly, for moderate variations of the cut (–), the normalized distribution exhibits only minor differences and remains consistent within uncertainties. This indicates that our baseline selection captures the relevant neutron-neutron correlation physics without either overly truncating compact configurations or introducing spurious distortions.
I.11 Distribution of the dihedral angle in 7H
To quantify the non-coplanarity of four-neutron configurations in 7H, we introduce the dihedral angle between the two planes spanned by each neutron pair and the c.m. of 7H. Using the neutron coordinates relative to the 7H c.m., we construct the normal vectors of two pair-c.m. planes for each pinhole configuration. By construction, corresponds to approximately coplanar arrangements of the two planes. As increases, the planes open up and approach an orthogonal configuration near . Because a plane normal is defined only up to a sign (depending on the orientation convention of the in-plane vectors used to construct it), the dihedral angles and are physically equivalent. We therefore fold the distribution and restrict the dihedral angle to without loss of information. Figure S5 shows the normalized probability density for different selections of the inter-pair opening angle . For the inclusive selection , is relatively broad, indicating that the sampled configurations span a wide range of torsional geometries, while the weight near the coplanar limit () is suppressed. For two selected intervals, the distributions for exhibit a pronounced peak at intermediate dihedral angles, -, whereas those for shift strength toward larger dihedral angles and peak near .
@@ I.11 Distribution of the dihedral angle in 7H
I.12 Internal angular and radial distributions from the four-neutron correlations in 7H
@@
Figure S6 provides additional information on the internal angular and radial distributions in the four-neutron correlation analysis and serves as a consistency check against the two-neutron correlations in 7H. Panels (a,e) compile the joint probability densities and for the two neutron-neutron pairs identified in each sampled four-neutron configuration. The bulk of the probability density is concentrated around the typical dineutron angular and radial scales extracted independently from the two-neutron correlations (dashed lines), indicating that the leading internal scales inferred from two-body observables are consistent with those obtained from the explicit four-body correlation. This agreement supports the interpretation of the observed bimodal angular pattern in 7H from the two-body correlation analysis: the two-body correlation function represents a statistical distribution over all neutron-neutron pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. Panels (b,f), obtained with the selection , are similar to (a,e) and contain the majority of the sampled configurations, indicating that the (extended) sector dominates the sampled four-body ensemble. Panels (d,h) correspond to a characteristic window around a maxima of the marginal distribution (). For the large- window (d,h), the radial distribution shows that and are comparable, and shifts toward smaller intra-pair opening angles, consistent with a near back-to-back arrangement of the two pair c.m. vectors at comparable radii and a more compact intra-pair geometry. In contrast, the small- window (c,g) shows a more asymmetric joint radial distribution, featuring enhanced weight near fm, alongside a nearly symmetric contribution at fm. The intra-pair opening-angle distribution shifts to larger angles and becomes noticeably broader, with enhanced support around , compared to the tighter concentration near in panel (d). The associated torsion- and dihedral-angle distributions are also weighted toward large angles in our calculation, indicating a strongly noncoplanar and interlaced four-neutron geometry with pronounced four-body correlations.